# The environment of TeV halo progenitors

### v1

@@ 2.4 Supernova remnant evolution
The SNR parameters are the SN energy, for which we used the fiducial value of $E_{\mathrm{SN}}=10^{51}\leavevmode\nobreak\ \mathrm{erg}$, and the ejecta mass $m_{\mathrm{ej}}$. It could be possible to add a dispersion on $E_{\mathrm{SN}}$, but observationally, this parameter is degenerate with the medium density, resulting in loose constraints on the parameter, so we kept it at its commonly assumed value. The ejecta mass is computed as the progenitor mass to which we subtracted the mass lost in winds and the remainder of the pulsar mass $m_{\mathrm{PSR}}=1.4\leavevmode\nobreak\ \mathrm{M}_{\odot}$:
@@
The model is based on the method given in Appendix A of Ptuskin & Zirakashvili (2005) for the computation of the SNR position $r$ as a function of time $t$. We recall the equations we use here, beginning by the equation for the SNR mass $M(r)=M_{\mathrm{ej}}+M_{\mathrm{sw}}(r)$ with $M_{\mathrm{sw}}(r)$ the swept-up mass until the shock position $r$:
@@
The SNR mass enters in the shock velocity $u_{\mathrm{SNR}}$ profile expression
$u_{\mathrm{SNR}}(r)=\frac{\gamma+1}{2}\left[\frac{2\alpha E_{\mathrm{SN}}}{M^{2}(r)r^{\alpha}}\underbrace{\int_{0}^{r}\mathrm{d}r^{\prime}r^{\prime\alpha-1}M(r^{\prime})}_{I(r)}\right]^{1/2}$ (25)
with $\gamma=5/3$ the adiabatic index for a monoatomic gas in the medium and $\alpha=6(\gamma-1)/(\gamma+1)$.
@@
The SNR speed profile can be solved analytically by assuming a density profile. We recall that the one we have follows:
@@
We solved Eq. 25 analytically in this density profile and Eq. 26 numerically to finally find the SNR’s age as a function of its position.
When and if the SNR reaches the bubble shell, it has swept the mass emitted by the progenitor star in the form of MS and RSG winds, but also some interstellar matter evaporated from the shell as a result of thermal conduction (Weaver et al. 1977). Yet, the shell made up of swept-up ISM material is much heavier than the SNR and effectively acts as a solid wall. As a result, the SNR collides and very rapidly loses most of its energy radiatively, as discussed both theoretically and observationally in Vink (2020) to merge with the shell. In our model framework, the SNR therefore has an impact on the bubble interior, but never reaches the ISM.
In case the SNR does not encounter the shell, it can merge with its surrounding medium at the position $r_{\mathrm{merge}}$ when it stops being a strong shock
@@
where $\beta=2$ is an arbitrary parameter that gives the minimal shock strength and $c_{\mathrm{s}}(r)=\sqrt{\gamma k_{\mathrm{B}}T(r)/\mu m_{\mathrm{p}}}$ the speed of sound in the medium.
@@ 3 Synthetic population of pulsars
In this section, we present the results of our model on the type of environment crossed by a moving pulsar as a function of time, up to pulsar ages of a few Myr. In the following, we will use the concept of ’parent environment’ to refer to the structured medium that surrounds a pulsar at birth. This immediate-vicinity environment is determined by the nature of the progenitor star (isolated or in a cluster), as well as by the properties of the supernova remnant resulting from the core-collapse explosion, whose expansion modifies the surrounding medium. We first introduce the properties of this parent environment in relation to our model parameters, before presenting the statistical implications for the medium experienced by pulsars along their trajectory.
@@ 3.1 Properties of the parent environment
Figure 1 shows the evolution of SNRs for the two extreme values of progenitor star mass in the range we considered, $8\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and $40\leavevmode\nobreak\ \mathrm{M}_{\odot}$, along with the corresponding wind-bubble structure at the time of supernova. In both cases, the expansion of the SNR is halted by its encounter with the bubble shell, at $3\leavevmode\nobreak\ \mathrm{kyr}$ and $52\leavevmode\nobreak\ \mathrm{kyr}$ respectively. This is due to the mass of the bounding shell being much higher than that of the expanding remnant, by at least two orders of magnitude (see Table 1). In the WBB case, a typical pulsar with a kick velocity of $350\leavevmode\nobreak\ \mathrm{km/s}$ escapes from the bubble shell and enters the ISM at $64\leavevmode\nobreak\ \mathrm{kyr}$ and $153\leavevmode\nobreak\ \mathrm{kyr}$, respectively.
@@
Figure 2 shows the evolution of SNRs for a progenitor star of mass $8\leavevmode\nobreak\ \mathrm{M}_{\odot}$ in a SB, for three different values of the cluster mass, $10^{3}\leavevmode\nobreak\ \mathrm{M}_{\odot}$, $10^{4}\leavevmode\nobreak\ \mathrm{M}_{\odot}$, and $10^{5}\leavevmode\nobreak\ \mathrm{M}_{\odot}$. As expected, the higher the cluster mass, the larger the SB radius because of the more powerful collective wind. In these examples, rather large SB are obtained because the chosen progenitor star has a long lifetime, during which the bubble could expand to large distances. Since the parameters of the explosion are the same, the trajectory changes because of the density structure inside the bubble cavity, which is primarily determined by the cluster luminosity and age, and by the exterior density. We see in all three cases that the SNR evolves from an ejecta-dominated phase to a Sedov-Taylor phase, before merging inside the SB without even reaching the bubble shell. When this happens, a typical pulsar with a kick velocity of $350\leavevmode\nobreak\ \mathrm{km/s}$ is still located inside the remnant.
@@
We notice that the SB radius in the most massive case of $10^{5}\leavevmode\nobreak\ \mathrm{M}_{\odot}$ is $320\leavevmode\nobreak\ \mathrm{pc}$, which is larger than the typical scale height of the gas disk of our Galaxy (Langer et al. 2014). A proper treatment would require to model the evolution of the superbubble in at least two dimensions with a gas density gradient, as well as a proper modelling of the stellar population in the cluster but this is beyond the scope of our study. In addition, because of the cluster IMF, this case is statistically rare, so we simply neglect these considerations.
@@
In the adopted model framework, the WBBs and SBs shells are the relevant boundaries marking the entry into the ISM, and not the forward shock of the SNR (as would be the case when not taking into account a modified circumstellar medium). Before a pulsar reaches the ISM, it can be either inside its parent SNR or inside the WBB or SB (owing to their small widths, we neglect the short time spent in the bounding shells of the bubbles). When the pulsar resides in the bubble medium, it is interesting to distinguish two different setups: one in which the SNR is still alive and expanding, and one in which it has dissolved in the WBB or SB interior. Yet, in our Monte-Carlo simulations and given the adopted pulsar kick velocities, the latter case appears to be more common. In the WBB case, a pulsar can escape a still-expanding SNR for a kick velocity higher than $1000\leavevmode\nobreak\ \mathrm{km/s}$, and only for a $\sim 40\leavevmode\nobreak\ \mathrm{M}_{\odot}$ progenitor. Both this progenitor mass and kick velocity are rare, hence there is a small probability of such a setup being found in nature. In the SB case, prospects are more interesting. A pulsar can escape a still-expanding SNR and be found in the SB interior for a cluster mass $10^{3}\leavevmode\nobreak\ \mathrm{M}_{\odot}$, which is the most favoured by the cluster IMF, but only if it has a kick velocity higher than $800\leavevmode\nobreak\ \mathrm{km/s}$, and a very massive progenitor of $40\leavevmode\nobreak\ \mathrm{M}_{\odot}$. However, the most probable situation for the SB case is that the SNR merges with the bubble cavity while the pulsar was still moving within. Observationally, such a setup could have relevant consequences, especially in the context of pulsar halos, since it implies the possibility to find a young or middle-aged pulsar in what may very likely be a turbulent medium (the interior of the bubble), while the parent SNR is not there anymore to be detected. Geminga may very well be in such a situation.
@@
Figure 1: Evolution of the SNR as a function of time in the WBB scenario, for two values of the stellar progenitor initial mass. For each case, we overplotted the positions of the wind-termination shock (dotted lines) and outer bubble radius (dashed lines) at the time of explosion. The orange line displays the trajectory of a pulsar with kick velocity $350\leavevmode\nobreak\ \mathrm{km/s}$.
@@
Figure 2: Evolution of the SNR as a function of time in the SB scenario, for a stellar progenitor of initial mass $8\leavevmode\nobreak\ \mathrm{M}_{\odot}$ and three different masses of the stellar cluster it belongs to. Graphical elements are similar to Fig. 1. The SNR curves turn flat when the remnant merges with the cavity interior and disappears.
@@ 3.2 Statistical description of a pulsar environment
From these mock populations, we computed the probability as a function of age that a pulsar is located in a particular medium among those two: the interior of the SNR and the interior of the WBB or SB. A null probability means that the pulsar has escaped into the ISM. We plotted the corresponding results in Fig. 3, using as pulsar kick velocity distribution that of Igoshev (2020) since it is based on the most complete pulsar sample. For comparison, we over-plotted as vertical colored lines the characteristic ages of the first four pulsar halos discovered by HAWC (B0656+14 and Geminga) and LHAASO (J0622+3749 and J0248+6021), and as a shaded band the classical values of $40-60\leavevmode\nobreak\ \mathrm{kyr}$ routinely found in the literature for the time at which a pulsar enters the ISM (e.g., Gaensler & Slane 2006; Amato & Recchia 2024).
@@
Our population synthesis actually predicts that pulsars will enter the ISM somewhat later than usually assumed: at an age of 60 kyr, more than $60\%$ of pulsars in the WBB case and $90\%$ in the SB case still reside in their parent environment. As expected due to the smaller size of the structure, pulsars remain inside WBBs for a shorter amount of time compared to SBs. In the latter case, the transition from pulsars being mostly in to pulsar being mostly out occurs at an age of $200-300$ kyr.
@@
Both the WBB and SB cases can be combined, with the proper weighing for the global proportion of isolated versus clustered massive stars (see Sect. 2.3), to make predictions for a Galactic population. This is illustrated in Fig. 4, which resembles strongly the SB case in Fig. 3 since most massive stars live in clusters. Overall, it can be seen that before $30\leavevmode\nobreak\ \mathrm{kyr}$, almost all pulsars are inside their parent environment, while all escaped after $5\leavevmode\nobreak\ \mathrm{Myr}$. Most pulsars escape into the ISM at $\sim 100-300$ kyr, with half of the Galactic population being out at $200\leavevmode\nobreak\ \mathrm{kyr}$. This is precisely the age range of the so-called middle-aged pulsars presumably involved in TeV halos.
@@
From our model and population synthesis, there is therefore no clear-cut interpretation of TeV halos in terms of the medium a pulsar lies in, since middle-aged pulsars have comparable chances of being inside their (possibly turbulent) parent environment or in the (a priori more quiescent) interstellar medium. Conversely, if we assume that the parent environment, provides enough turbulence for efficient trapping of ultrarelativistic electron-positron pairs, and that this is the main formation channel for TeV halos, our model predicts a strongly decreasing occurrence rate of TeV halos as pulsars get older than 100 kyr. At the characteristic age of Geminga, only a third of pulsars would then be able to form halos. Our model also provides the complementary information that about half of these halo-forming pulsars would still be located inside their parent SNR (see Fig. 3). This is qualitatively consistent with the proposal put forward in Martin et al. (2022) that pulsar halos may not be so common around middle-aged pulsars.
@@
Finally, in Fig. 4, we illustrated the impact of the kick velocity distribution adopted in the Monte-Carlo sampling. We see that there is a clear impact on the position of the transition from predominantly in or out of the parent environment, and to a lesser extent on its shape. As a consequence, the quantitative statements made above or below should be treated with some caution.
@@
Figure 3: Probability for a pulsar to be found inside its parent remnant or bubble environment as a function of time. Red and blue lines correspond to the SB and WBB cases, respectively. The grey-shaded band corresponds to values typically found in the literature for the time at which a pulsar exits into the ISM, thus showing that they are significantly underestimated. The four vertical lines correspond to the characteristic ages of the confirmed TeV halos.
@@ 4.1 Sample construction
We applied our model to a sample of real pulsars in order to derive constraints on their environments, in a statistical sense. From the ATNF catalogue (Manchester et al. 2005), we extracted a subset of pulsar halo candidates having properties relatively similar to Geminga, the prototypical pulsar halo.
We considered pulsars with luminosity to distance ratio $\dot{E}/d\geq 10^{33}\leavevmode\nobreak\ \mathrm{erg/s/kpc}$ and characteristic age $40\leavevmode\nobreak\ \mathrm{kyr}<\tau_{\mathrm{age}}<1\leavevmode\nobreak\ \mathrm{Myr}$. The first criterion aims at jointly maximizing the power and angular size of the halo emission, while the second is meant to discard too young systems that are more likely to be classical PWNe. We added a criterion to restrict our sample to relatively nearby objects. For a TeV halo to be detected as an extended source with current experiments, an upper limit needs to be set on the distance of the system. Assuming a typical halo size of $r_{\mathrm{halo}}\sim 30\leavevmode\nobreak\ \mathrm{pc}$ and considering that it would be detected as extended with HAWC and LHAASO if spanning at least 3 point-spread functions of $\alpha_{\mathrm{PSF}}\sim 0.3\leavevmode\nobreak\ \mathrm{deg}$, we derive an upper limit on the halo visibility at distance $d_{\mathrm{max}}=r_{\mathrm{halo}}/\tan(\alpha_{\mathrm{PSF}})=2\leavevmode\nobreak\ \mathrm{kpc}$.
We complemented the sample with the four well-established pulsar TeV halos (J0633+1746, B0656+14, J0622+3749 and J0248+6021) and three other objects not fulfilling the above criteria but mentioned in the literature as good halo candidates, mainly on the basis of gamma-ray observations (B0540+23, J0633+0632, and J0631+1036; see Celli & Peron 2024; Khokhriakova et al. 2024).
Finally, our set of real pulsars was divided into those having estimated transverse velocity and those that do not, because the former category allows us to make a finer treatment and to derive richer constraints on their environments. The full sample is listed in Table 2.

### v2

@@ 2.4 Supernova remnant evolution
The SNR parameters are the SN energy, for which we used the fiducial value of $E_{\mathrm{SN}}=10^{51}\penalty 10000\ \mathrm{erg}$, and the ejecta mass $m_{\mathrm{ej}}$. It could be possible to add a dispersion on $E_{\mathrm{SN}}$, but observationally this parameter is degenerate with the medium density, resulting in loose constraints on the parameter. Therefore, we kept it at its commonly assumed value. The ejecta mass was computed as the progenitor mass minus the mass lost in winds and the remaining pulsar mass $m_{\mathrm{PSR}}=1.4\penalty 10000\ \mathrm{M}_{\odot}$:
@@
The model is based on the method given in Appendix A of Ptuskin and Zirakashvili (2005) for the computation of the SNR position $r$ as a function of time $t$. We recall the equations used here, beginning with the equation for the SNR mass $M(r)=M_{\mathrm{ej}}+M_{\mathrm{sw}}(r)$, with $M_{\mathrm{sw}}(r)$ representing the swept-up mass until the shock position $r$:
@@
The SNR mass enters in the shock velocity $u_{\mathrm{SNR}}$ profile expression:
$u_{\mathrm{SNR}}(r)=\frac{\gamma+1}{2}\left[\frac{2\alpha E_{\mathrm{SN}}}{M^{2}(r)r^{\alpha}}\underbrace{\int_{0}^{r}\mathrm{d}r^{\prime}r^{\prime\alpha-1}M(r^{\prime})}_{I(r)}\right]^{1/2},$ (26)
with $\gamma=5/3$ being the adiabatic index for a monatomic gas in the medium and $\alpha=6(\gamma-1)/(\gamma+1)$.
@@
The SNR speed profile can be solved analytically by assuming a density profile. We recall that the density profile we use follows:
@@
We solved Eq. 26 analytically in this density profile and Eq. 27 numerically to finally find the SNR’s age as a function of its position. When and if the SNR reaches the bubble shell, it sweeps the mass emitted by the progenitor star in the form of MS and RSG winds, as well as some interstellar matter evaporated from the shell as a result of thermal conduction (Weaver et al. 1977). Yet, the shell made up of swept-up ISM material is much heavier than the SNR and effectively acts as a solid wall. As a result, the SNR collides and very rapidly loses most of its energy radiatively, as discussed both theoretically and observationally in Vink (2020), to merge with the shell. In our model framework, the SNR therefore has an impact on the bubble interior but never reaches the ISM. In case the SNR does not encounter the shell, it can merge with its surrounding medium at the position $r_{\mathrm{merge}}$ when it stops being a strong shock:
@@
where $\beta=2$ (Cioffi et al. 1988) is an arbitrary parameter that gives the minimal shock strength and $c_{\mathrm{s}}(r)=\sqrt{\gamma k_{\mathrm{B}}T(r)/\mu m_{\mathrm{p}}}$ is the speed of sound in the medium.
@@ 3 Synthetic population of pulsars
In this section, we present the results of our model on the type of environment crossed by a moving pulsar as a function of time, up to pulsar ages of a few megayears. In the following, we use the concept of ’parent environment’ to refer to the structured medium that surrounds a pulsar at birth. This immediate-vicinity environment is determined by the nature of the progenitor star (isolated or in a cluster), as well as by the properties of the SNR resulting from the core-collapse explosion, whose expansion modifies the surrounding medium. We first introduce the properties of this parent environment in relation to our model parameters, before presenting the statistical implications for the medium experienced by pulsars along their trajectory.
@@ 3.1 Properties of the parent environment
[…]
@@
Figure 2 shows the evolution of SNRs for a progenitor star of mass $8\penalty 10000\ \mathrm{M}_{\odot}$ in a SB, for three different values of the cluster mass, $10^{3}\penalty 10000\ \mathrm{M}_{\odot}$, $10^{4}\penalty 10000\ \mathrm{M}_{\odot}$, and $10^{5}\penalty 10000\ \mathrm{M}_{\odot}$. As expected, the higher the cluster mass, the larger the SB radius because of the more powerful collective wind. In these examples, rather large SB are obtained because the chosen progenitor star has a long lifetime, during which the bubble could expand to large distances. Since the parameters of the explosion are the same, the trajectory changes because of the density structure inside the bubble cavity, which is primarily determined by the cluster luminosity and age, and by the exterior density. We see in all three cases that the SNR evolves from an ejecta-dominated phase to a Sedov-Taylor phase, before merging inside the SB without even reaching the bubble shell. When this happens, a typical pulsar with a kick velocity of $350\penalty 10000\ \mathrm{km/s}$ remains inside the remnant.
@@
We notice that the SB radius in the most massive case of $10^{5}\penalty 10000\ \mathrm{M}_{\odot}$ is $320\penalty 10000\ \mathrm{pc}$, which is larger than the typical scale height of the gas disc of our Galaxy (Langer et al. 2014). A proper treatment would require us to model the evolution of the SB in at least 2D with a gas density gradient, as well as a proper modelling of the stellar population in the cluster but this is beyond the scope of our study. In addition, because of the cluster IMF, this case is statistically rare, so we simply neglect these considerations.
@@
In the adopted model framework, the WBB and SB shells are the relevant boundaries marking the entry into the ISM, and not the forward shock of the SNR (as would be the case when not taking into account a modified circumstellar medium). Before a pulsar reaches the ISM, it can be either inside its parent SNR, the WBB, or the SB (owing to their small widths, we neglect the short time spent in the bounding shells of the bubbles). When the pulsar resides in the bubble medium, it is interesting to distinguish two different setups: one in which the SNR is still alive and expanding, and one in which it has dissolved in the WBB or SB interior. Yet, in our Monte-Carlo simulations and given the adopted pulsar kick velocities, the latter case appears to be more common. Given the range of progenitor masses, in the WBB case, the SNR always reaches the bubble shell before the pulsar, namely the pulsar can either be found inside the SNR or in the ISM. In the SB case, the prospects are more interesting. A pulsar can escape a still-expanding SNR and be found in the SB interior for a cluster mass $10^{3}\penalty 10000\ \mathrm{M}_{\odot}$ – the most favoured by the cluster IMF – but only if it has a kick velocity typically exceeding $800\penalty 10000\ \mathrm{km/s}$ and a massive progenitor of $20\penalty 10000\ \mathrm{M}_{\odot}$. However, the most probable situation for the SB case is that the SNR merges with the bubble cavity while the pulsar was still moving within. Observationally, such a setup could have relevant consequences, particularly in the context of pulsar haloes, since it implies the possibility of finding a young or middle-aged pulsar in what is likely a turbulent medium (the interior of the bubble), even though the parent SNR can no longer be detected. Geminga may very well be in such a situation.
@@
Figure 1: Evolution of the SNR as a function of time in the WBB scenario, for two values of the stellar progenitor initial mass. For each case, we overplotted the positions of the wind-termination shock (dotted lines, on top of each other, see Tab. 1) and outer bubble radius (dashed lines) at the time of explosion. The purple line indicates the trajectory of a pulsar with kick velocity $350\penalty 10000\ \mathrm{km/s}$.
@@
Figure 2: Evolution of the SNR as a function of time in the SB scenario, for a stellar progenitor of initial mass $8\penalty 10000\ \mathrm{M}_{\odot}$ and three different masses of the stellar cluster it belongs to. Graphical elements are similar to Fig. 1. The SNR curves turn flat when the remnant merges with the cavity interior and disappears.
@@ 3.2 Statistical description of a pulsar environment
From these mock populations, we computed the probability as a function of age that a pulsar is located in a particular medium among those two: the interior of the SNR and the interior of the WBB or SB. A null probability means that the pulsar has escaped into the ISM. We plotted the corresponding results in Fig. 3, using as pulsar kick velocity distribution that of Igoshev (2020) since it is based on the most complete pulsar sample. For comparison, we overplotted as vertical coloured lines the characteristic ages of the first four pulsar haloes discovered by HAWC (B0656+14 and Geminga) and LHAASO (J0622+3749 and J0248+6021), and as a shaded band the classical values of $40-60\penalty 10000\ \mathrm{kyr}$ routinely found in the literature for the time at which a pulsar enters the ISM (e.g. Gaensler and Slane 2006; Amato and Recchia 2024).
@@
Our population synthesis actually predicts that pulsars will enter the ISM somewhat later than usually assumed: at an age of 60 kyr, more than $52\%$ of pulsars in the WBB case and all pulsars in the SB case still reside in their parent environment. As expected due to the smaller size of the structure, pulsars remain inside WBBs for a shorter amount of time compared to SBs. In the latter case, the transition from pulsars being mostly in to pulsar being mostly out occurs at an age of 200500200-500 kyr.
@@
Both the WBB and SB cases can be combined, with the proper weighing for the global proportion of isolated versus clustered massive stars (see Sect. 2.3), to make predictions for a Galactic population. This is illustrated in Fig. 4, which resembles strongly the SB case in Fig. 3 since most massive stars live in clusters. Overall, it can be seen that before $30\penalty 10000\ \mathrm{kyr}$, almost all pulsars are inside their parent environment, while all escaped after $6\penalty 10000\ \mathrm{Myr}$. Most pulsars escape into the ISM at $\sim 100-1000$ kyr, with half of the Galactic population being out at $300\penalty 10000\ \mathrm{kyr}$. This is precisely the age range of the so-called middle-aged pulsars presumably involved in TeV haloes.
@@
From our model and population synthesis, there is therefore no clear-cut interpretation of TeV haloes in terms of the medium a pulsar lies in, since middle-aged pulsars have comparable chances of being inside their (possibly turbulent) parent environment or in the (a priori more quiescent) ISM. Conversely, if we assume that the parent environment, provides enough turbulence for efficient trapping of ultrarelativistic electron-positron pairs, and that this is the main formation channel for TeV haloes, our model predicts a strongly decreasing occurrence rate of TeV haloes as pulsars get older than 300 kyr. At the characteristic age of Geminga, half of the pulsars would then be able to form haloes. Our model also provides the complementary information that about half of these halo-forming pulsars would still be located inside their parent SNR (see Fig. 3). This is qualitatively consistent with the proposal put forward in Martin et al. (2022) that pulsar haloes may not be so common around middle-aged pulsars.
As can be see in Fig. 2 in the SB case, the SNR can merge before reaching the bubble shell. A pulsar can be observed without a SNR surrounding it but still being inside the SB. Increasing the required shock strength before merger from $\beta=2$ to $\beta=5$ makes the SNR merge with the interior of the bubble marginally earlier, slightly increasing the time the pulsar spends inside the SB without probing the downstream of a SNR.
@@
Finally, in Fig. 4, we illustrate the impact of the kick velocity distribution adopted in the Monte-Carlo sampling. We see a clear impact on the position of the transition – whether predominantly in or out of the parent environment – and, to a lesser extent, on its shape. As a consequence, the quantitative statements made above or below should be treated with some caution.
@@
Figure 3: Probability for a pulsar to be found inside its parent remnant or bubble environment as a function of time. The red and blue lines correspond to the SB and WBB cases, respectively. The grey-shaded band corresponds to values typically found in the literature for the time at which a pulsar exits into the ISM, thus showing that they are significantly underestimated. The four vertical lines correspond to the characteristic ages of the confirmed TeV haloes.
@@ 4.1 Sample construction
We applied our model to a sample of real pulsars in order to derive constraints on their environments, in a statistical sense. From the ATNF catalogue (Manchester et al. 2005), we extracted a subset of pulsar halo candidates with properties that are relatively similar to Geminga, the prototypical pulsar halo.
We considered pulsars with luminosity to distance ratio $\dot{E}/d\geq 10^{33}\penalty 10000\ \mathrm{erg/s/kpc}$ and characteristic age $40\penalty 10000\ \mathrm{kyr}<\tau_{\mathrm{age}}<1\penalty 10000\ \mathrm{Myr}$. The first criterion aims at jointly maximising the power and angular size of the halo emission, while the second is meant to discard that are too young systems and are more likely to be classical PWNe. We added a criterion to restrict our sample to relatively nearby objects. For a TeV halo to be detected as an extended source with current experiments, an upper limit needs to be set on the distance of the system. Assuming a typical halo size of $r_{\mathrm{halo}}\sim 30\penalty 10000\ \mathrm{pc}$ and considering that it would be detected as extended with HAWC and LHAASO if spanning at least three point-spread functions of $\alpha_{\mathrm{PSF}}\sim 0.3\penalty 10000\ \mathrm{deg}$, we derived an upper limit on the halo visibility at distance $d_{\mathrm{max}}=r_{\mathrm{halo}}/\tan(\alpha_{\mathrm{PSF}})=2\penalty 10000\ \mathrm{kpc}$. We complemented the sample with the four well-established pulsar TeV haloes (J0633+1746, B0656+14, J0622+3749 and J0248+6021) and three other objects not fulfilling the above criteria but mentioned in the literature as good halo candidates, mainly on the basis of gamma-ray observations (B0540+23, J0633+0632, and J0631+1036; see Celli and Peron 2024; Khokhriakova et al. 2024). Finally, our set of real pulsars was divided into those having estimated transverse velocity and those without, as the former category allows us to make a finer treatment and to derive richer constraints on their environments. The full sample is listed in Table LABEL:table:pulsar_properties.